OpenStudy (anonymous):

how do you Match the power with its value if it has a negitive exponent? ex: 8^-2

OpenStudy (saifoo.khan):

\[\Large 8^{-2} \to \frac{1}{8^2} \to \frac{1}{64}\]

OpenStudy (anonymous):

A negative exponent brings the term that the exponent is being applied, into it's reciprocal. 8^-2 = 1/8^2 = 1/64